position vectors of points A, B and C are 3i-2j-k, 2i+3j-4k, -i+j+2k and 4i+5j+ak respectively. if points A, B, C and D lie on a plane, find the value of a.

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
AB=B-A=2i+3j-4k- (3i-2j-k) = -i+5j+3k
BC=C-B=-3i-2j+6k
CD=5i+4j+(a-2)k
z= AB*BC=36 i + 15j -13 k

z.cd= 180+60-13a+26=0

a=266/13